# Learning Encodings by Maximizing State Distinguishability: Variational Quantum Error Correction

### v1

@@ B.2 Connection to Fidelity after Recovery
The central goal of our paper is to learn resource-efficient encoding operations $U_{\mathrm{enc}}(\Theta)$ that protect quantum information against given noise channels $\mathcal{N}$. This ability is quantified by the worst-case distinguishability loss $\overline{\mathcal{D}}(\mathcal{N};\Theta)$. However, to perform successful quantum error correction, one also has to construct and perform a recovery operation $U_{\mathrm{rec}}(\Phi)$. How to learn a recovery operation we elaborate superficially in the main part of this paper, and discuss in more detail in Appendix C. In the end, the measure of interest is how well the recovered (and decoded) state aligns with the original state, which we quantify by the worst-case fidelity loss ¯$\overline{\mathcal{F}}(\mathcal{N};\Theta,\Phi)$. In the empirical sections of this paper, we demonstrate for various noise models and setups that an encoding with lower distinguishability loss allows for a recovery operation with higher fidelity, i.e. lower fidelity loss. While it is not essential for the main statements of our paper, we think it is still valuable to derive a rigorous relation between these two quantities. Therefore, we state and prove the following theorem, which was also informally stated in the main part:
@@
In simpler words, Theorem 1 states that decreasing the distinguishability loss for an encoding operation guarantees a lower fidelity loss, i.e. an improved fidelity after recovery. In practice, this bound is rarely tight, as this relation holds already purely due to the existence of a trivial recovery operation $U_{\mathrm{rec}}=\mathbb{I}$, as used in the proof. Therefore, usually it should be possible to find much better recovery operations than guaranteed by this Theorem 1. The experimental results in this paper, generated with the procedure for learning recovery operations described in Appendix C, confirm this assumption. However, we also empirically observe that a lower distinguishability loss allows for a higher-fidelity recovery operation.
@@
¯$\overline{\mathcal{F}}(\mathcal{N})\geq\overline{\mathcal{D}}(\mathcal{N})^{2}.$ (86)
@@
Complementary to the above, Theorem 2 states that it is not only sufficient but also necessary to achieve a low distinguishability loss, to allow for a high-fidelity recovery operation. However, especially for small $\overline{\mathcal{D}}$, this bound is close to $1$, i.e. almost trivial.
@@ B.3 Code Distance and Approximate Error Correction
which by definition allows it to detect $d-1=w$ errors and correct for
@@
$[[n,k,d]]$ (92)
to classify the error correction capabilities of a code. In this work, we consider QEC codes with $n\leq 10$, $k\leq 2$, and $d\leq 4$, but in the main part, we also sketch potential improvements that could allow scaling to larger code parameters.
@@
$[[n,k,d^{*}]].$ (95)
In the case of Definition 1, it would also be possible to alternatively use the average-case code distinguishability loss $\mathcal{D}_{\mathcal{S}}$ instead of the worst-case formulation $\overline{\mathcal{D}}_{\mathcal{S}}$. We observed that in all cases the strongest signal $\overline{\mathcal{D}}_{\mathcal{S}}(\mathcal{N}_{p})$ is exhibited for $p=0.5$, i.e. for empirical evaluations we focus on this setup.
@@
$[[n,k,d^{*}_{\varepsilon}]].$ (98)
$1$ fault $2$ faults $3$ faults $4$ faults
@@
Table 2: Evaluation of potential (approximative) code distance following Definitions 1 and 2 for different QEC and VarQEC codes. We perform evaluations for $1$ to $4$ faults, corresponding to the respective error weights in Eq. 88. The number column denotes the according count of possible error configurations from Eq. 89 that are evaluated. A checkmark ✓ for detect indicates that Eq. 93 is satisfied for this code and error weight, for (✓) at least the weaker Eq. 96 holds with $\varepsilon=0.002$. For ✗ neither holds, which also indicates the potential (approximative) code distance of the respective code.
Based on these definitions, we report empirical evaluations of several standard QEC codes, and two instances of VarQEC codes in Table 2. In particular, we determine the potential code distance following Definition 1 for standard (non-approximative) codes with known code distance, and the potential approximate code distance for codes learned on depolarizing and amplitude damping noise.
@@
¯$\overline{\mathcal{F}}(\mathcal{N}_{p})=1$ (100)
for all noise channels $\mathcal{N}_{p}(\rho)=(1-p)\cdot\rho+p\cdot E\rho E^{\dagger}$ where $E$ has a weight of $w\leq\lfloor\frac{d^{*}-1}{2}\rfloor$. We note that the results from Theorems 1 and 2 are not applicable here, as these noise channels do not satisfy the stated requirements.
Alternatively to this framework of perfect error correction, there is also the notion of approximate error correction [84, 10], which relaxes Eq. 100 to
¯$\overline{\mathcal{F}}(\mathcal{N}_{p})\geq 1-\delta$ (101)
for some small positive value $\delta$. For completeness, we notice that both perfect and approximate error correction conditions can also be stated in terms of the Knill-Laflamme conditions [51], but the notion using the fidelity is easier to interpret for our work. By construction, the proposed VarQEC procedure does not explicitly enforce the ability to recover information under specific error terms. Rather, the distinguishability loss is used to automatically focus on specific error terms, to minimize the overall loss of information. Consequently, approximate error correction is the more fitting framework to analyze these codes. Similar to above, we hypothesize that a VarQEC code with parameters [[$n,k,d^{*}_{\varepsilon}$]] allows for the correction of up to $\lfloor\frac{d^{*}_{\varepsilon}-1}{2}\rfloor$ with a high fidelity following Eq. 101, where $\delta$ correlates with $\varepsilon$. Interestingly, the [[$5,1$]]-VarQEC code that is analyzed in Fig. 7(a) practically performs equivalent to the [[$5,1,3$]]-QEC code, i.e. we can assume $\delta$ being very close to 00, or at least below numerical accuracy of the evaluation procedure. Consequently, it is reasonable to interpret this instance of a VarQEC-code as being capable of performing practically exact error correction.
@@ B.4 Approximation with Two-Designs
To allow for an efficient evaluation of both the average-case distinguishability loss $\mathcal{D}$ and worst-case distinguishability loss $\overline{\mathcal{D}}$, we introduced two-design approximations $\mathcal{D}_{\mathcal{S}}$ and $\overline{\mathcal{D}}_{\mathcal{S}}$, respectively. For the exact definitions, please refer to the main part or Eqs. 49, 50, 52 and 53 in the appendix. The non-approximative instances are based on integrating or maximizing over pairs of Haar-random pure $k$-qubit states [89], respectively. Empirically, these operations have to be mimicked by averaging or maximizing over a finite set of $H$ states following a Haar-random distribution.
In Fig. 9, we analyze how many such states are actually necessary for $k=1$ and $k=2$ to get a stable approximation to the ground-truth. In the main parts of this paper, we base our evaluations on $H=1000$ Haar-random states, which should guarantee a deviation from the ground truth of at most $0.001$ in the $k=1$ case and $0.01$ for $k=2$. To save computational resources, in the appendix we reduce this to $H=200$, which still should limit the deviations to about $0.005$.
@@
Figure 9: Empirical analysis of how many Haar-random states are necessary to get a stable approximation of the ground-truth loss, which integrates or maximizes over an infinite number of states. The loss fluctuations denote the maximal difference between distinguishability loss values over $10$ instances of the same encoding, with different random seeds for generating the Haar-random states. The data points denote the median, while the error bars show the smallest/largest fluctuation over $5$ different encoding circuits. The plots refer to depolarizing noise with $p=0.1$, but similar results can be obtained for other noise models.
@@
[…]
On the right side of Eq. 102, $\rho$ are Haar-random quantum states. In principle, the distinguishability loss contains degrees of up to order $4$, i.e. one would require a $4$-design to guarantee this equivalence. However, we evaluate below that in practice, using a two-design, which is much easier to construct and evaluate, allows for a sufficiently accurate approximation.
@@
This spherical two-design originating from a single qubit is also visualized in Fig. 2(c). For two data qubit, i.e. $k=2$, we use a formulation with $12$ product states and $4$ entangled states, reading as
@@
In Fig. 10, we analyze how well we can approximate the ground-truth distinguishability loss with the respective two-design formulations. By the definitions, it must, in principle, hold
@@
This is because the state pair creating the largest loss of trace distance within the two-design must trivially be reflected in the consideration of all state pairs. For the average-case formulation, no such direct relation exists. Empirically, we quantify the approximation quality as $\mathcal{D}/\mathcal{D}_{\mathcal{S}}$ and $\overline{\mathcal{D}}/\overline{\mathcal{D}}_{\mathcal{S}}$, respectively. For both the $k=1$ setup in Fig. 10(a) and the $k=2$ setup in Fig. 10(b), we can see that while there is some bias in average-case approximation, the variance of results is not too large. In contrast, for the worst-case approximation, the bias is smaller, but the variance is larger. The points above $1.0$, which in principle violate Eq. 108, can be explained by the fact that the baseline is only estimated with a finite number of Haar-random states. Moreover, we do not need to care too much about the worst-case approximation quality, as for evaluating this loss we usually directly resort to $\overline{\mathcal{D}}$ instead of $\overline{\mathcal{D}}_{\mathcal{S}}$. For training, we employ the average-case approximation $\mathcal{D}_{\mathcal{S}}$, but as it is used as a trainable loss function, the approximate proportionality to the ground truth $\mathcal{D}$ is sufficient.
@@
In Fig. 11 we exemplarily consider an amplitude damping channel with damping parameter $\gamma=0.1$. With this channel, low-energy states like |$\ket{0}$ are less or not affected at all, while high-energy states like |$\ket{1}$ are disproportionally strongly affected. We reflect this in setting the weights as
@@
(a) Approximation with standard 2-design.
(b) Approximation with weighted 2-design.
Figure 11: Empirical analysis on how a weighted two-design can reduce the approximation bias, compared to a standard two-design on an amplitude damping channel with $\gamma=0.1$. Hereby, we evaluate $100$ random encoding circuit instances for each setup, and estimate the ground truth with $H=200$ Haar-random states. the weights are set as summarized in Eq. 111.
@@ Appendix C Learning Recovery Operations with Fidelity Loss
To set the state, let us start with the definition of an encoded state $\rho_{L}$ encoded with the potentially parameterized encoding operation $U_{\mathrm{enc}}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(\Theta)}$ as described in Eq. 51. After a noise channel $\mathcal{N}$ acts on this state, the task is now to find a recovery operation $U_{\mathrm{rec}}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(\Phi)}$, which reverts the effect of $\mathcal{N}$ as much as possible. Especially for the multi-recovery setting, i.e. multiple consecutive recovery operations, this requires the use of additional recovery qubits initialized to a known state, usually |$\ket{0}^{\otimes r}$. Most established QEC procedures use this register for measuring stabilizers and extracting syndromes [35]. Subsequently, classically controlled correction operations are applied to the encoded states. While it should be possible to adapt the presented VarQEC routine within this framework, it more naturally aligns with the concept of measurement-free or autonomous QEC [54, 105, 41]. After applying the recovery operation, the register is traced out, which in practice could be done by re-initializing to |$\ket{0}^{\otimes r}$ or supplying new qubits for the next recovery cycle. Considering the single-recovery setup, the resulting state $\hat{\rho}_{L}$ reads as
$\hat{\rho}_{L}=\mathrm{Tr}_{r}\left(U_{\text{rec}}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(\Phi)}\left(\mathcal{N}({\rho}_{L})\otimes\ket{0}\bra{0}^{\otimes r}\right)U_{\text{rec}}^{\dagger}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(\Phi)}\right)$ (112)
@@
where $\rho$ should be close to $\hat{\rho}$ for a successful QEC cycle. This closeness can be quantified by the fidelity $F(\rho,\hat{\rho})$. This is the underlying idea of the QVECTOR approach, which uses variational ansätze for both the encoding and recovery operation, and defines the average-case fidelity loss as
@@
¯$\overline{\mathcal{F}}(\mathcal{N}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5};\Theta,\Phi})=1-\min_{\rho}F(\rho,\hat{\rho}).$ (115)
@@
$F(p\otimes\ket{0}^{n-k},U^{\dagger}_{\mathrm{enc}}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(\Theta)}\hat{\rho}_{L}U_{\mathrm{enc}}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(\Theta)}).$ (116)
@@
The fidelity loss trivially converges to 00 if the fidelity between original and recovered state is close to $1$, i.e. one can formulate and optimize the QVECTOR objective as
@@
This procedure is also sketched in Fig. 12. The end-to-end QVECTOR approach proves sophisticated enough for noise models like phase and amplitude damping with small damping parameters [46]. However, it struggles with more complex noise models and larger noise rates. In the experimental section of this paper, we demonstrate that our two-step procedure succeeds on a variety of noise models and does so with a much shallower circuit, which is beneficial for fault-resilience and fault-tolerance.
@@
Figure 12: The adapted QVECTOR pipeline [46] for training recovery operations $U_{\mathrm{rec}}(\Phi)$, under the premise that the encoding operations $U_{\mathrm{enc}}(\Theta)$ have already been trained with the distinguishability loss. The average-case fidelity loss $\mathcal{F}_{\mathcal{S}}$ is evaluated on states from a two-design $\mathcal{S}$ and is minimized w.r.t. the free parameters of the recovery operation. We note that at the end of the circuits it is not guaranteed that the state $\hat{\rho}$ on the $k$ data qubits is completely disentangled from the $n-k$ ancilla qubits, but for the single-recovery scenario, this can be neglected.

### v2

@@ B.2 Connection to Fidelity after Recovery
The central goal of our paper is to learn resource-efficient encoding operations $U_{\mathrm{enc}}(\Theta)$ that protect quantum information against given noise channels $\mathcal{N}$. This ability is quantified by the worst-case distinguishability loss $\overline{\mathcal{D}}(\mathcal{N};\Theta)$. However, to perform successful quantum error correction, one also has to construct and perform a recovery operation $U_{\mathrm{rec}}(\Phi)$. How to learn a recovery operation we elaborate superficially in the main part of this paper, and discuss in more detail in Appendix˜C. In the end, the measure of interest is how well the recovered (and decoded) state aligns with the original state, which we quantify by the worst-case fidelity loss $\overline{\mathcal{F}}(\mathcal{N};\Theta,\Phi)$. In the empirical sections of this paper, we demonstrate for various noise models and setups that an encoding with lower distinguishability loss allows for a recovery operation with higher fidelity, i.e. lower fidelity loss. While it is not essential for the main statements of our paper, we think it is still valuable to derive a rigorous relation between these two quantities. Therefore, we state and prove the following theorem, which was also informally stated in the main part:
@@
In simpler words, Theorem˜1 states that decreasing the distinguishability loss for an encoding operation guarantees a lower fidelity loss, i.e. an improved fidelity after recovery. In practice, this bound is rarely tight, as this relation holds already purely due to the existence of a trivial recovery operation $U_{\mathrm{rec}}=\mathbb{I}$, as used in the proof. Therefore, usually it should be possible to find much better recovery operations than guaranteed by this Theorem˜1. The experimental results in this paper, generated with the procedure for learning recovery operations described in Appendix˜C, confirm this assumption. However, we also empirically observe that a lower distinguishability loss allows for a higher-fidelity recovery operation.
@@
$\overline{\mathcal{F}}(\mathcal{N})\geq\overline{\mathcal{D}}(\mathcal{N})^{2}.$ (90)
@@
Complementary to the above, Theorem˜2 states that it is not only sufficient but also necessary to achieve a low distinguishability loss, to allow for a high-fidelity recovery operation. However, especially for small $\overline{\mathcal{D}}$, this bound is close to $1$, i.e. almost trivial.
@@ B.3 Code Distance and Approximate Error Correction
which by definition allows it to detect d1=wd-1=w errors and correct for
@@
$((n,2^{k},d)),$ (96)
or $[[n,k,d]]$ for stabilizer codes, to classify the error correction capabilities of a code. In this work, we consider QEC codes with $n\leq 10$, $k\leq 2$, and $d\leq 4$, but in the main part, we also sketch potential improvements that could allow scaling to larger code parameters.
@@
$((n,2^{k},d^{*})).$ (99)
In the case of Definition˜1, it would also be possible to alternatively use the average-case code distinguishability loss $\mathcal{D}_{\mathcal{S}}$ instead of the worst-case formulation $\overline{\mathcal{D}}_{\mathcal{S}}$. We observed that in all cases the strongest signal $\overline{\mathcal{D}}_{\mathcal{S}}(\mathcal{N}_{p})$ is exhibited for $p=0.5$, i.e. for empirical evaluations we focus on this setup.
@@
$((n,2^{k},d^{*}_{\varepsilon})).$ (102)
11 fault 22 faults 33 faults 44 faults
@@
Table 2: Evaluation of potential (approximative) code distance following Definitions˜1 and 2 for different QEC and VarQEC codes. We perform evaluations for 11 to 44 faults, corresponding to the respective error weights in equation˜92. The number column denotes the according count of possible error configurations from equation˜93 that are evaluated. A checkmark ✓ for detect indicates that equation˜97 is satisfied for this code and error weight, for (✓) at least the weaker equation˜100 holds with $\varepsilon=0.002$. For ✗ neither holds, which also indicates the potential (approximative) code distance of the respective code.
Based on these definitions, we report empirical evaluations of several standard QEC codes, and two instances of VarQEC codes in table˜2. In particular, we determine the potential code distance following Definition˜1 for standard (non-approximative) codes with known code distance, and the potential approximate code distance for codes learned on depolarizing and amplitude damping noise.
@@
$\overline{\mathcal{F}}(\mathcal{N}_{p})=1$ (104)
for all noise channels $\mathcal{N}_{p}(\rho)=(1-p)\cdot\rho+p\cdot E\rho E^{\dagger}$ where $E$ has a weight of $w\leq\lfloor\frac{d^{*}-1}{2}\rfloor$. We note that the results from Theorems˜1 and 2 are not applicable here, as these noise channels do not satisfy the stated requirements.
Alternatively to this framework of perfect error correction, there is also the notion of approximate error correction [90, 10], which relaxes equation˜104 to
$\overline{\mathcal{F}}(\mathcal{N}_{p})\geq 1-\delta$ (105)
for some small positive value $\delta$. For completeness, we notice that both perfect and approximate error correction conditions can also be stated in terms of the Knill-Laflamme conditions [55], but the notion using the fidelity is easier to interpret for our work. By construction, the proposed VarQEC procedure does not explicitly enforce the ability to recover information under specific error terms. Rather, the distinguishability loss is used to automatically focus on specific error terms, to minimize the overall loss of information. Consequently, approximate error correction is the more fitting framework to analyze these codes. Similar to above, we hypothesize that a VarQEC code with parameters $((n,2^{k},d^{*}_{\varepsilon}))$ allows for the correction of up to $\lfloor\frac{d^{*}_{\varepsilon}-1}{2}\rfloor$ with a high fidelity following equation˜105, where $\delta$ correlates with $\varepsilon$. Interestingly, the $((5,2))$ VarQEC code that is analyzed in figure˜7(a) practically performs equivalent to the $[[5,1,3]]$ QEC code, i.e. we can assume $\delta$ being very close to 0, or at least below numerical accuracy of the evaluation procedure. Consequently, it is reasonable to interpret this instance of a VarQEC-code as being capable of performing practically exact error correction.
@@ B.4 Approximation with Two-Designs
To allow for an efficient evaluation of both the average-case distinguishability loss $\mathcal{D}$ and worst-case distinguishability loss $\overline{\mathcal{D}}$, we introduced two-design approximations $\mathcal{D}_{\mathcal{S}}$ and $\overline{\mathcal{D}}_{\mathcal{S}}$, respectively. For the exact definitions, please refer to the main part or equations˜53, 54, 56 and 57 in the appendix. The non-approximative instances are based on integrating or maximizing over pairs of Haar-random pure $k$-qubit states [96], respectively. Empirically, these operations have to be mimicked by averaging or maximizing over a finite set of $H$ states following a Haar-random distribution.
In figure˜9, we analyze how many such states are actually necessary for $k=1$ and $k=2$ to get a stable approximation to the ground-truth. In the main parts of this paper, we base our evaluations on $H=1000$ Haar-random states, which should guarantee a deviation from the ground truth of at most $0.001$ in the $k=1$ case and $0.01$ for $k=2$. To save computational resources, in the appendix we reduce this to $H=200$, which still should limit the deviations to about $0.005$.
@@
Figure 9: Empirical analysis of how many Haar-random states are necessary to get a stable approximation of the ground-truth loss, which integrates or maximizes over an infinite number of states. The loss fluctuations denote the maximal difference between distinguishability loss values over $10$ instances of the same encoding, with different random seeds for generating the Haar-random states. The data points denote the median, while the error bars show the smallest/largest fluctuation over 55 different encoding circuits. The plots refer to depolarizing noise with $p=0.1$, but similar results can be obtained for other noise models.
@@
$\frac{1}{S}\sum_{\rho_{k}\in\mathcal{X}}P_{t}(\rho_{k})=\int_{\rho}P_{t}(\rho)\,d\nu(\rho)$ (106)
On the right side of equation˜106, $\rho$ are Haar-random quantum states. In principle, the distinguishability loss contains degrees of up to order $4$, i.e. one would require a $4$-design to guarantee this equivalence. However, we evaluate below that in practice, using a two-design, which is much easier to construct and evaluate, allows for a sufficiently accurate approximation.
@@
This spherical two-design originating from a single qubit is also visualized in figure˜2(c). For two data qubit, i.e. $k=2$, we use a formulation with $12$ product states and $4$ entangled states, reading as
@@
In figure˜10, we analyze how well we can approximate the ground-truth distinguishability loss with the respective two-design formulations. By the definitions, it must, in principle, hold
@@
This is because the state pair creating the largest loss of trace distance within the two-design must trivially be reflected in the consideration of all state pairs. For the average-case formulation, no such direct relation exists. Empirically, we quantify the approximation quality as $\mathcal{D}/\mathcal{D}_{\mathcal{S}}$ and $\overline{\mathcal{D}}/\overline{\mathcal{D}}_{\mathcal{S}}$, respectively. For both the $k=1$ setup in figure˜10(a) and the $k=2$ setup in figure˜10(b), we can see that while there is some bias in average-case approximation, the variance of results is not too large. In contrast, for the worst-case approximation, the bias is smaller, but the variance is larger. The points above $1.0$, which in principle violate equation˜112, can be explained by the fact that the baseline is only estimated with a finite number of Haar-random states. Moreover, we do not need to care too much about the worst-case approximation quality, as for evaluating this loss we usually directly resort to $\overline{\mathcal{D}}$ instead of $\overline{\mathcal{D}}_{\mathcal{S}}$. For training, we employ the average-case approximation $\mathcal{D}_{\mathcal{S}}$, but as it is used as a trainable loss function, the approximate proportionality to the ground truth $\mathcal{D}$ is sufficient.
@@
(a) Approximation with standard 2-design.
(b) Approximation with weighted 2-design.
Figure 11: Empirical analysis on how a weighted two-design can reduce the approximation bias, compared to a standard two-design on an amplitude damping channel with $\gamma=0.1$. Hereby, we evaluate $100$ random encoding circuit instances for each setup, and estimate the ground truth with $H=200$ Haar-random states. the weights are set as summarized in equation˜115.
@@
[…]
@@ Appendix C Learning Recovery Operations with Fidelity Loss
[…]
@@
$\mathcal{F}(\mathcal{N}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5};\Theta,\Phi})=1-\int_{\rho}F(\rho,\hat{\rho})\,d\nu(\rho),$ (118)
@@
[…]
@@
[…]
@@
The fidelity loss trivially converges to 0 if the fidelity between original and recovered state is close to $1$, i.e. one can formulate and optimize the QVECTOR objective as
@@
This procedure is also sketched in figure˜12. The end-to-end QVECTOR approach proves sophisticated enough for noise models like phase and amplitude damping with small damping parameters [50]. However, it struggles with more complex noise models and larger noise rates. In the experimental section of this paper, we demonstrate that our two-step procedure succeeds on a variety of noise models and does so with a much shallower circuit, which is beneficial for fault-resilience and fault-tolerance.
@@
Figure 12: The adapted QVECTOR pipeline [50] for training recovery operations $U_{\mathrm{rec}}(\Phi)$, under the premise that the encoding operations $U_{\mathrm{enc}}(\Theta)$ have already been trained with the distinguishability loss. The average-case fidelity loss $\mathcal{F}_{\mathcal{S}}$ is evaluated on states from a two-design $\mathcal{S}$ and is minimized w.r.t. the free parameters of the recovery operation. We note that at the end of the circuits it is not guaranteed that the state $\hat{\rho}$ on the $k$ data qubits is completely disentangled from the nkn-k ancilla qubits, but for the single-recovery scenario, this can be neglected.